-2<6x+7/8< or = 2
need a step by step

Answer by stanbon(75887)

-2<6x+7/8< or = 2
-----------------------
-2 < (6x+(7/8)) <= 2
--------------------
Multiply thru by 8 to get rid of the fraction:
-16 < 48x+7 <= 16
Subtract 7 along the line:
-23 < 48x <= 9
Divide thru by 48 to solve for "x":
-(23/48) < x <= (9/48)
===========================
Cheers,
Stan H.